OpenStudy (anonymous):

HELP!! n an 1 4 2 12 3 28 4 60 Which recursive equation represents the pattern? an = an – 1 + 2n an = an – 1 + 2(n + 1) an = an – 1 + 2n + 1 an = an – 1 + 2n + 1

OpenStudy (anonymous):

All you need to do is to plug in both a and an and see which one is equal.

OpenStudy (anonymous):

yeah, but what is the equation?

OpenStudy (mathstudent55):

12 - 4 = 8 28 - 12 = 16 60 - 28 = 32 Why are the third and fourth choices equal?

OpenStudy (anonymous):

You find which one is the right one by trying every equation till you get both sides equal.

OpenStudy (mathstudent55):

Are some n's and (n-1)'s subscripts and exponents?

OpenStudy (anonymous):

I'm still a lot confused on what the heck to do...

OpenStudy (anonymous):

@mathstudent55 What do I subtract 60 by?

OpenStudy (anonymous):

Grr.. Im so confused!! :/

OpenStudy (mathstudent55):

n an 1 4 2 12 4 + 8 4 + 2^3 3 28 12 + 16 12 + 2^4 4 60 28 + 32 28 + 2^5 Do you see a pattern? Each term, an, is obtained by adding the previous term to 2 to the power n + 1

OpenStudy (mathstudent55):

The way you wrote the choices makes no sense. There must be some exponents and subscripts there. You wrote everything as if it were simple numbers.

OpenStudy (mathstudent55):